5. (10 pts.) A 30-kV stopping potential is required to stop a Li particle (m -1.1510 q +2e, e 1.602x10 1 c). (a) What is the initial energy of the particle in electron volts? (b) What is the initial speed of the particle?
Answer #1

here,

stopping potential , V = 30 KV = 30000 V

mass of particle , m = 1.15 * 10^-26 kg

charge , q = 2 e

a)

using work energy theorm

the initial energy of the particle , KE = V * q

KE = 30000 * 2 * e

KE = 6 * 10^4 eV

b)

let the initial speed of particle be u

KE = 0.5 * m * u^2

6 * 10^4 * e = 0.5 * m * u^2

6 * 10^4 * 1.6 * 10^-19 = 0.5 * 1.15 * 10^-26 * u^2

solving for u

u = 1.3 * 10^6 m/s

the initial speed of particle is 1.3 * 10^6 m/s